OpenStudy (anonymous):

(a + b)(a^2 - ab + b^2) im sorry but i need help on this one too

OpenStudy (anonymous):

I think that is a sum of cubes

OpenStudy (ageta):

(a+b)(a^2−ab+b^2) =(a+b)(a^2+−ab+b^2) =(a)(a^2)+(a)(−ab)+(a)(b^2)+(b)(a^2)+(b)(−ab)+(b)(b^2) =a^3−a^2b+ab^2+a^2b−ab^2+b^3 =a^3+b^3

OpenStudy (anonymous):

Ageta is correct

OpenStudy (anonymous):

(a+b)(a2−ab+b2) =(a+b)(a2+−ab+b2) =(a)(a2)+(a)(−ab)+(a)(b2)+(b)(a2)+(b)(−ab)+(b)(b2) =a3−a2b+ab2+a2b−ab2+b3 =a3+b3

OpenStudy (ageta):

thank you @Mathaintcool

OpenStudy (anonymous):

For those type of questions use Mathpapa, great algebra calculator shows you how to do the work as well

OpenStudy (anonymous):

http://www.mathpapa.com/algebra-calculator.html

OpenStudy (anonymous):

thank you :)

OpenStudy (ageta):

your welcome

OpenStudy (anonymous):

Any more questions? :P

OpenStudy (anonymous):

you know you can go to https://mathway.com/

OpenStudy (anonymous):

give you correct answers in seconds

OpenStudy (anonymous):

Mathways doesnt show how to do it Ameliana

OpenStudy (anonymous):

Unless you donate ;)

OpenStudy (anonymous):

donate to wat

OpenStudy (anonymous):

Mathways, if you dont donate you dont get to see the work on how to solve the problems

OpenStudy (anonymous):

http://www.acalculator.com/math-calculators.html